A. 1/6 B. 5/8 C. 3/8 D. 5/6 Explanation:Advertisement P(at least one graduate) = 1 – P(no graduates) = 1 – ⁶C₃/¹⁰C₃ = 1 – (6 * 5 * 4)/(10 * 9 * 8) = 5/6 Advertisement